OpenStudy (anonymous):

Can you show me how to solve the next five in this series of fraction 23/24, 11/12, 21/24. 5/6, 19/24, 3/4

OpenStudy (amistre64):

the bottom is alternating between 24 and 12: the next number is has a 24... make em all the same denominator and see how the top is changing...

OpenStudy (anonymous):

Wouldn't you just divide the first and second to find the rate of change?

OpenStudy (anonymous):

I still don't understand

OpenStudy (amistre64):

to see the pattern here; turn everything into fractions with the SAME denominator.

OpenStudy (amistre64):

the tops will do this if you do: 23,22,21,20,19,18,17,16,15,14,13,12.... The next one in the series is: 17/24 if I kept track of it :)

OpenStudy (amistre64):

23/24 = 23/24 11/12 = 22/24 21/24 = 21/24 5/6 = 20/24 see the pattern??